A number is four less than four times another number. If the product of the two numbers is 63. Find the two numbers?
Answer by ankor@dixie-net.com(22740)

Two numbers: x & y
:
"A number is four less than four times another number."
x = 4y - 4
:
If the product of the two numbers is 63.
x*y = 63
:
Find the two numbers?
:
Replace x with (4y-4)
(4y-4)*y = 63
4y^2 - 4y = 63
4y^2 - 4y - 63 = 0
Factor
(2y-9)(2y+7) = 0
two solutions
2y = 9
y = 4.5
and
2y = -7
y = -3.5
:
Find x when y = 4.5
x = 4(4.5) -4
x = 18 - 4
x = 14 (14 * 4.5 = 63)
:
Find x when y = -3.5
x = 4(-3.5) - 4
x = -14 - 4
x = -18 (-18 * -3.5 = 63)
:
Two solutions 14, 4.5 and -18, -3.5